"please solve these problem"
the sum of elisa's age and elma's age is 20. elisa's age one year from now will ten times elma's age one year ago. how old is each now.
thank you.


Answer by stanbon(75887)

the sum of elisa's age and elma's age is 20. elisa's age one year from now will ten times elma's age one year ago. how old is each now.
--------------------------------------------------
Let elisa be x yrs old
Let elma be y years old
----------------------------------
NOW Equation: x + y = 20
1 yr from now Eq; (x+1) = 10(y-1)
---------------------------------------
Rearrange:
x + y = 20
x -10y=-11
--------------
Subtract the 2nd from the first:
11y = 31
y = (31/11) = 2.8181 years old (elma/s age NOW)
-------
Since x+y = 20, x = 17.1818... (elisa's age NOW)
====================================================
Cheers,
Stan H.